OpenStudy (anonymous):

A rock dropped from a high platform is moving at 24 m/s downward when it strikes the ground. Ignore air resistance. How fast was the rock moving when it had fallen only one-fourth of the distance to the ground?

OpenStudy (ybarrap):

Which of these equations do you think we should use? http://en.wikipedia.org/wiki/SUVAT_equations#SUVAT_equations Hint: We have s, g and u.

OpenStudy (anonymous):

s=ut + 1/2 *a * t ^2

OpenStudy (anonymous):

Would be my guess

OpenStudy (ybarrap):

There are a few steps, but not hard: If you use v=u-gt You can find out how long it took to hit the ground Then you can use s=(1/2)(u+v)t to find out how far the rock actually fell. Take (1/4) of the value of s Then find out how long it would take to fall this distance using (s/4)=ut-(1/2)gt^2 Then use s/2=(1/2)(u+v)t to find out how fast it was going.

OpenStudy (ybarrap):

I meant s/4 where I have s/2 above, sorry about that.

OpenStudy (ybarrap):

This is because we are interested in 1/4 the distance not 1/2 the distance.

OpenStudy (anonymous):

Alright, so the first part is 2.45 seconds, I think. 24 = 0 - 9.8t 24 = -9.8t -2.45 = t (but time obviously can't be negative)! Look good so far?

OpenStudy (ybarrap):

But velocity is negative because downward is negative. Also note that s is negative, because s is going down. Looks good so far.

OpenStudy (ybarrap):

Note that g was also negative

OpenStudy (anonymous):

Ah, okay Next part... 0.5 * 24 * -2.45 = s = -29.4 -29.4 * 1/4 = -7.35

OpenStudy (anonymous):

-7.35 = 0 * -2.45 - 1/2 (0+24) * -2.45

OpenStudy (anonymous):

Wait... I messed something up :/

OpenStudy (anonymous):

I'm confused again...

OpenStudy (ybarrap):

You are here: You have s/4 and u=0, g, now solve for t: (s/4)=ut-(1/2)gt^2

OpenStudy (anonymous):

Oh, Okay! So it should look like this: -7.35 = 0 * t - 1/2 (0+24) * t Right?

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

That should be 9.8 * t^2 at the end

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

So the next part is -7.35 = -1/2 * 9.8 * t^2

OpenStudy (ybarrap):

yes, so $$ t=\sqrt{\cfrac{2\times 7.35}{9.8}}~seconds $$

OpenStudy (anonymous):

So t = 1.22

OpenStudy (ybarrap):

I'll let you check calculations, that's the easy part. I'll check formulas.

OpenStudy (anonymous):

Sounds good :)

OpenStudy (anonymous):

s/4 = (1/2) (u+v) t solving for v

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

7.35 = 1/2 * v * 1.22

OpenStudy (ybarrap):

Solve for v and you are done!

OpenStudy (anonymous):

So v = 12! I think haha

OpenStudy (anonymous):

That was a long one. Thank you so much for your help :)

OpenStudy (ybarrap):

No problem, I did check that last calculation and got the same. But be sure to double check all your numbers. Formulas are okay. You're welcome.

OpenStudy (anonymous):

Is there any chance you could offer me some advice on two other problems? You won't have to walk me through those, I just want an idea of how I might solve them.

OpenStudy (anonymous):

Here is the first one: A roller-coaster car starts at a height of 90 m at Point A and is moving at a speed of 20 m/s on a later hill at Point C. Assuming friction does not affect the motion of the car, what is its height at Point C? And here is the second one: A 0.145 kg baseball is thrown at launch angle of 30° and strikes the ground at 18 m/s. How fast would it be moving when it reaches the ground if its launch angle were 45°? Ignore air resistance.

OpenStudy (anonymous):

@wolfe8 Is there any chance you can point me in the right direction with these two questions? You don't need to help much, but I would appreciate any help :)!

OpenStudy (wolfe8):

For the new questions, can you make new posts for them?